# Subleading correction to the Asian options volatility in the Black-Scholes model

### v2

@@ 2. The asymptotic distribution of $\frac{1}{t}A_{t}^{(\mu)}$
(28) $f_{HW}(a,\tau):=\frac{1}{2\pi t}e^{-\frac{1}{2}\mu^{2}t}a^{\mu-1}\int_{0}^{\infty}\rho^{\mu}G(\rho)e^{-\frac{1}{\tau}H(\rho,a)}\frac{d\rho}{\rho}\,.$
@@
The application of Laplace asymptotic methods to the integral (28) gives the following result.
Proposition 3.
[Proposition 6 in [21]] We have the $\tau\to 0$ asymptotics
(31) $f_{HW}(a,\tau)=\frac{1}{\sqrt{2\pi t}}g(a,\mu)e^{-\frac{1}{t}J(a)}\frac{da}{a}(1+O(\tau))$
where
Substituting the result (31) into (27) gives an approximation $f_{0}(a,\tau)$ for the density of the time integral of the gBM. The error of this approximation is bounded by the following result.
@@
(34) $|f(a,\tau)-f_{HW}(a,\tau)|\leq e^{-\frac{1}{2}\mu^{2}\tau}a^{\mu-1}\int_{0}^{\infty}\rho^{\mu}e^{-\frac{1+a^{2}\rho^{2}}{2a\tau}}|\theta(\rho/\tau,\tau)-\frac{1}{2\pi\tau}e^{-\frac{1}{t}[F(\rho)-\frac{\pi^{2}}{2}]}G(\rho)|\frac{d\rho}{\rho}$
$\hskip 56.9055pt\leq\frac{1}{2\pi\tau}e^{-\frac{1}{2}\mu^{2}t}a^{\mu-1}\int_{0}^{\infty}\rho^{\mu}G(\rho)e^{-\frac{1}{\tau}H(\rho,a)}|\vartheta(\rho,\tau)|\frac{d\rho}{\rho}\leq\frac{1}{70}\tau f_{HW}(a,\tau)$
@@
(38) $\frac{1-\frac{1}{70}\tau}{1+\frac{1}{70}\tau}f(a,\tau)\leq f_{0}(a,\tau)\leq\frac{1+\frac{1}{70}\tau}{1-\frac{1}{70}\tau}f(a,\tau)$
@@
The function $f_{HW}(a,\tau)$ depends on two functions $J(a)$ and $g(a,\mu)$. The properties of the function $J(a)$ were studied in Sec. 4.1 of [21] where it was shown that it is simply related to the rate function $J_{BS}(k)$ appearing in the short-maturity asymptotics of Asian options, as
@@
$g(a,\mu)$ $=$ $e^{\mu\log a+(\mu-1)\log\rho_{*}(a)}G(\rho_{*})\frac{1}{\sqrt{H^{\prime\prime}(\rho_{*})}}$
@@
(42) $c_{2}=-\frac{3}{80}(\mu+1)+\frac{57}{1,400}$
(43) $c_{3}=\frac{1}{350}(\mu+1)-\frac{1}{875}\,.$
@@ 3.3. Including a non-zero interest rate $r$
$L$ $=$ $\frac{1}{\tau}J(e^{x})+(\mu+1)e^{x}J^{\prime}(e^{x})-\frac{3}{2}\log\tau+\alpha_{0}(e^{x})+O(\tau)$
@@
::=:= $\frac{1}{\tau}J(e^{x})-\frac{3}{2}\log\tau+\tilde{\alpha}_{0}(e^{x})+O(\tau)\,.$

### v3

@@ 2. The asymptotic distribution of $\frac{1}{t}A_{t}^{(\mu)}$
(28) $f_{HW}(a,\tau):=\frac{1}{2\pi\tau}e^{-\frac{1}{2}\mu^{2}\tau}a^{\mu-1}\int_{0}^{\infty}\rho^{\mu}G(\rho)e^{-\frac{1}{\tau}H(\rho,a)}\frac{d\rho}{\rho}\,.$
@@
The error of the approximation (27) is bounded by the following result.
@@
(32) $|f(a,\tau)-f_{HW}(a,\tau)|\leq e^{-\frac{1}{2}\mu^{2}\tau}a^{\mu-1}\int_{0}^{\infty}\rho^{\mu}e^{-\frac{1+a^{2}\rho^{2}}{2a\tau}}|\theta(\rho/\tau,\tau)-\frac{1}{2\pi\tau}e^{-\frac{1}{\tau}[F(\rho)-\frac{\pi^{2}}{2}]}G(\rho)|\frac{d\rho}{\rho}$
$\hskip 56.9055pt\leq\frac{1}{2\pi\tau}e^{-\frac{1}{2}\mu^{2}\tau}a^{\mu-1}\int_{0}^{\infty}\rho^{\mu}G(\rho)e^{-\frac{1}{\tau}H(\rho,a)}|\vartheta(\rho,\tau)|\frac{d\rho}{\rho}\leq\frac{1}{70}\tau f_{HW}(a,\tau)$
@@
(36) $\frac{1-\frac{1}{70}\tau}{1+\frac{1}{70}\tau}f(a,\tau)\leq f_{0}(a,\tau)\leq\frac{1+\frac{1}{70}\tau}{1-\frac{1}{70}\tau}f(a,\tau)\,.$
@@
In this paper we are interested in the small-$\tau$ expansion of the integral (28). The application of Laplace asymptotic methods to this integral gives the more explicit result.
Proposition 4.
[Proposition 6 in [21]] We have the $\tau\to 0$ asymptotics
(37) $f_{HW}(a,\tau)=\frac{1}{\sqrt{2\pi\tau}}g(a,\mu)e^{-\frac{1}{\tau}J(a)}(1+O(\tau))$
where
The leading asymptotics of the function $f_{HW}(a,\tau)$ in (37) depends on two functions $J(a)$ and $g(a,\mu)$. The properties of the function $J(a)$ were studied in Sec. 4.1 of [21] where it was shown that it is simply related to the rate function $J_{BS}(k)$ appearing in the short-maturity asymptotics of Asian options, as
@@
(40) $g(a,\mu)$ $=e^{\mu\log a+(\mu-1)\log\rho_{*}(a)}G(\rho_{*})\frac{1}{\sqrt{H^{\prime\prime}(\rho_{*})}}$
@@
$c_{2}=-\frac{3}{80}(\mu+1)+\frac{57}{1,400}$
$c_{3}=\frac{1}{350}(\mu+1)-\frac{1}{875}\,.$
@@ 3.3. Including a non-zero interest rate $r$
(71) $L$ $=\frac{1}{\tau}J(e^{x})+(\mu+1)e^{x}J^{\prime}(e^{x})-\frac{3}{2}\log\tau+\alpha_{0}(e^{x})+O(\tau)$
@@
$:=\frac{1}{\tau}J(e^{x})-\frac{3}{2}\log\tau+\tilde{\alpha}_{0}(e^{x})+O(\tau)\,.$